OpenStudy (anonymous):

Can someone please help me with this? Part a.) Write the equation of the line that has a slope of -2 and a y-intercept of -6. (5 points) Part b.) Write the equation of the line that passes through the point (6, -2) and has a slope of -2/3. (5 points)

OpenStudy (anonymous):

am from pakstain .and its too tough

jigglypuff314 (jigglypuff314):

were you taught point slope form or slope-intercept form?

OpenStudy (anonymous):

We were are freshman year, but i'm a junior now, it's be en a while since iv'e done this stuff.

jigglypuff314 (jigglypuff314):

for part a, you can use slope-intercept form which is: y = mx + b when m = slope and b = y-intercept so when the line has a slope of -2 and a y-intercept of -6, plug m = -2 and b = -6 into the slope-intercept form.

OpenStudy (anonymous):

So, y=-2(x)+(-6)

jigglypuff314 (jigglypuff314):

yep, that would be your first equation :)

jigglypuff314 (jigglypuff314):

now for the second equation, you can use point-slope form, which is: y - (y1) = m(x - (x1)) when given a point (x1, y1) and the slope = m so when given the point (6, -2) and a slope of -2/3 plug in x1 = 6, y1 = -2 , and m = (-2/3)

OpenStudy (anonymous):

y-2= -2/3(x-6)?

jigglypuff314 (jigglypuff314):

know that y - (-2) would be y + 2

OpenStudy (anonymous):

Okay, I think I got it, is that all I have to do?

jigglypuff314 (jigglypuff314):

yup :)

OpenStudy (anonymous):

Awesome! Thanks(: